OpenStudy (anonymous):

Please show your work and describe how to find an equation for the line with x-int (2, 0) and y-int (0,6). Thank you.

OpenStudy (anonymous):

slope = 6/-2 =-3 , ===> y=mx+b ==> y=-3x+6

OpenStudy (anonymous):

Why did the 2 become a negative?

OpenStudy (anonymous):

slope = (y2-y1)/(x2-x1) = (6-0)/(0-2) = 6/(-2) =-3

OpenStudy (anonymous):

that is by using the two given points.

OpenStudy (anonymous):

Thanks!